# From Tower to Spire: Adding the Speech Modality to a Text-Only LLM

### v1

@@ 4 Experiments
We evaluate our models across three tasks: ASR, MT, and ST. First, we present our results for ASR (§4.1), confirming the competitive performance of Spire in the speech domain. We then present MT results (§4.2), demonstrating that the speech performance does not come at the expense of the original model’s MT performance. Finally, we present results for ST (§4.3) to investigate model performance on a task that requires both ASR and MT capabilities.
Our Models
Across experiments, we compare all models shown in Table 4. These models vary in which base model they were trained on, whether they underwent speech-centric CPT, and on what set of instructions IT was performed. Apart from Spire models, we report a few ablation studies in the IT stage where:
i) no CPT was performed (TowerFull);
ii) no data from TowerBlocks was seen during IT (SpireNoBlocks), and
iii) pseudo-labeled ST data and FLEURS were omitted (SpireNoPseudo).
When reporting results, we highlight toplines when possible.
| Model | Base Model | CPT | IT | |||
| Speech | Text | Speech | Pseudo | Text | ||
| TowerFull | TowerBase-7B | ✗ | ✗ | ✓ | ✓ | ✓ |
| SpireBase | SpireBase | ✓ | ✓ | ✗ | ✗ | ✗ |
| SpireFull | SpireBase | ✓ | ✓ | ✓ | ✓ | ✓ |
@@
| SpireNoPseudo | SpireBase | ✓ | ✓ | ✓ | ✗ | ✓ |
Evaluation Setup
Across models and tasks, we perform inference with greedy decoding with a maximum of 256 generated tokens. For the Tower and Spire models, we decode with vllm. However, since vllm does not support all of our baselines, we use alternative libraries (namely, transformers) where necessary. Unless specified otherwise, we use zero-shot prompts for all models and tasks.
4.1 ASR Experiments
Datasets and Metrics
We evaluate ASR performance across multiple test sets, in order to cover a variety of recording styles: Librispeech (LS) test-clean and test-other Panayotov et al. (2015), FLEURS Conneau et al. (2023), and VoxPopuli.999For CPT models, LS is an in-domain evaluation because its training set is part of MLS. We report the Word Error Rate (WER) between the hypotheses and gold transcripts, after Whisper normalization (Radford et al., 2023).
Baselines
We report results for the following models:
Whisper (Radford et al., 2023) is an encoder-decoder transformer model trained on over 5 million hours of labeled data, it performs multilingual ASR and to-English ST. We report results for the 74M parameter Whisper-base and the 1.5B parameter Whisper-large-v3 versions.
SeamlessM4T (Barrault et al., 2023) is an encoder-decoder transformer trained on 406K hours of speech that performs ASR, ST and MT across 100 languages. We report results for the 2.3B parameter SeamlessM4T-large-v2 version of the model.
Spirit-LM Nguyen et al. (2025) is the most similar work to ours. It is a decoder-only model, trained from Llama-2 on 307B tokens of text, 458K hours of unlabeled speech, and 111K hours of labeled speech. Unfortunately, despite the availability of their inference code, we were unable to reproduce its reported performance on speech tasks.
HuBERT-large+CTC is a CTC-based ASR model trained using the same speech representation model we use for DSU generation, and using the same ASR data from the IT stage (Section 3.3.1). This model allows us to compare our IT-only TowerFull model against a model which has access to continuous speech representations.101010The hyperparameters for this ASR model are described in Appendix B.
| LibriSpeech | FLEURS | VoxPopuli | ||
| Clean | Other | |||
| Baselines | ||||
[…]
| HuBERT-large+CTC | 4.3 | 7.6 | 11.4 | 14.7 |
| Spirit-LM | 6.0* | 11.0* | - | - |
| SeamlessM4T | 2.6 | 4.9 | 8.1 | 7.5 |
| Whisper-base | 5.0 | 11.9 | 12.1 | 9.8 |
| Whisper-large-v3 | 1.8 | 3.7 | 5.8 | 9.2 |
| Our models | ||||
[…]
| TowerFull | 9.5 | 13.8 | 14.3 | 40.7 |
| SpireNoBlocks | 4.1 | 7.4 | 10.4 | 15.8 |
| SpireNoPseudo | 3.9 | 7.3 | 11.1 | 14.3 |
| SpireBase | 28.9 | 56.3 | 11.0 | 13.7 |
| SpireFull | 4.2 | 7.1 | 10.7 | 15.8 |
*We were unable to reproduce Spirit-LM’s ASR performance; therefore, we report their self-reported LS results using ten-shot prompts
@@ Results
Our results are presented in Table 4. SpireFull’s performance demonstrates that performing both the CPT and IT stages is an effective strategy to give speech capabilities to a text LLM. Notably, SpireFull outperforms the HuBERT-large+CTC baseline on three out of four datasets—an impressive result given that the CTC model has a helpful inductive bias and access to continuous features, both of which SpireFull lacks.
The performance gap between SpireFull and TowerFull (5.3 points in LS test-clean) demonstrates that combining CPT and IT is more effective than using IT alone. We further observe that while TowerFull obtains better results than SpireBase, it performs worse than HuBERT-large+CTC, showing that the CPT stage is crucial in outperforming a model that has access to continuous features. Additionally, the minimal difference between SpireNoBlocks and SpireFull in the IT stage suggests that incorporating textual tasks does not negatively impact ASR performance. For SpireBase, it is surprising that FLEURS and VoxPopuli results were somewhat strong zero-shot settings, given that non-instruction-tuned models often struggle to work in out-of-domain without in-context learning examples.111111We also tried prompting SpireBase with few-shot examples, but the results were significantly worse. This may be due to the length of the DSU sequences, which likely led to in-context examples that were too long for the model to handle effectively.
Finally, although SpireFull cannot match the performance of SeamlessM4T or Whisper-large-v3, both of which were trained on far more speech data, it does exceed the performance of Whisper-base on both sections of LS and FLEURS. It also outperforms Spirit-LM on LS, which is notable because both models are derived from Llama-2 and make use of HuBERT DSU tokens.
4.2 MT Experiments
@@
| TowerFull | 88.57 | 39.4 | 88.17 | 41.7 |
| SpireNoBlocks | 82.98 | 34.2 | 85.93 | 36.1 |
| SpireNoPseudo | 88.40 | 38.9 | 88.22 | 42.0 |
| SpireBase | 87.41 | 37.4 | 87.97 | 41.4 |
| SpireFull | 88.54 | 39.3 | 88.21 | 41.8 |
@@
| TowerFull | 82.63 | 39.2 | 84.55 | 34.5 | 85.39 | 37.2 | 84.65 | 45.2 | 82.41 | 35.6 | 79.68 | 25.9 |
| SpireNoBlocks | 67.97 | 24.4 | 73.86 | 26.6 | 77.80 | 29.6 | 73.24 | 28.7 | 78.09 | 29.1 | 73.01 | 17.6 |
| SpireNoPseudo | 82.18 | 38.5 | 84.31 | 34.7 | 85.31 | 37.6 | 85.04 | 45.5 | 82.56 | 36.2 | 79.91 | 26.2 |
| SpireBase | 79.88 | 34.7 | 83.04 | 33.7 | 83.85 | 32.4 | 83.19 | 40.5 | 80.20 | 32.4 | 78.65 | 23.1 |
| SpireFull | 82.50 | 39.5 | 84.60 | 34.9 | 85.37 | 37.3 | 85.24 | 45.2 | 82.58 | 36.4 | 79.92 | 26.3 |
Having demonstrated that our CPT and IT approaches work well for ASR, we now turn to MT. The key question is whether Spire can maintain Tower’s strong performance on MT, despite its speech-centric CPT. We report performance for translation-related tasks in Appendix C.
Datasets and Metrics
We use two datasets for MT: FLORES-200 Team et al. (2024), which covers all of our models’ languages, and the WMT23 test set (Kocmi et al., 2023), which covers en{de, ru, zh}. We report COMET-22 Rei et al. (2022a) and spBLEU121212nrefs:1|case:mixed|eff:no|tok:flores200| smooth:exp|version:2.5.1 Papineni et al. (2002) scores via the SacreBLEU toolkit Post (2018).
Baselines
We compare the Spire models against the text-to-text translation performance of SeamlessM4T. Additionally, we report the performance of Towerbase-7b and Towerinstruct-7b as toplines.
Results
Our results showcase that even after the speech-centric CPT and mixed speech and text IT stage, the Spire models retain Tower’s performance on both FLORES (Table 5) and WMT23 (Table 6). This indicates that neither CPT nor IT on speech data negatively impact the model’s ability to perform MT. This is true for both CPT-only models, where SpireBase achieves performance comparable to Towerbase on both datasets; and for IT models, where SpireFull and TowerFull both perform slightly better than Towerinstruct on enxx, which is possibly an artifact of the large number of multi-turn enxx ST instructions in their IT set. Notably, our strongest Spire model also outperforms SeamlessM4T by both metrics on all WMT23 language pairs, and for both enxx and xxen on FLORES.
4.3 ST Experiments
| FLEURS | CoVoST-2 | |||||||||||
| Direct | Self-Cascade | Gold | Direct | Self-Cascade | Gold | |||||||
| C22 | spB | C22 | spB | C22 | spB | C22 | spB | C22 | spB | C22 | spB | |
| Baselines | ||||||||||||
@@ 4.3 ST Experiments
| SeamlessM4T | 84.63 | 33.7 | 75.45 | 22.4 | 86.79 | 38.7 | 84.79 | 36.8 | 72.36 | 19.4 | 86.55 | 39.0 |
| Our Models | ||||||||||||
[…]
| TowerFull | 79.10 | 26.1 | 83.42 | 31.9 | 88.27 | 39.2 | 71.52 | 20.1 | 74.17 | 25.8 | 87.14 | 38.5 |
| SpireNoBlocks | 81.11 | 27.1 | 79.46 | 28.9 | 82.44 | 34.0 | 74.02 | 23.2 | 68.09 | 22.8 | 69.31 | 26.8 |
| SpireNoPseudo | 62.80 | 11.7 | 83.79 | 32.2 | 88.10 | 38.7 | 59.88 | 6.8 | 78.15 | 29.7 | 87.10 | 39.0 |
| SpireFull | 81.33 | 27.1 | 85.21 | 33.7 | 88.36 | 39.2 | 74.25 | 23.2 | 78.78 | 30.0 | 87.11 | 38.5 |
@@
We evaluate ST performance on both direct ST and self-cascaded ST,131313We also tried inference with the multi-turn prompt format shown in Table 1, but results were similar to a self-cascade. Reporting the self-cascade enables comparison to models that do not support the multi-turn format, i.e. SeamlessM4T. in which each model transcribes the audio before translating its own output to the target language (i.e., ASR followed by MT). To assess the impact of ASR error propagation, we also report MT results given gold transcriptions.
@@ Results
Our ST results on FLEURS and CoVoST-2 are presented in Table 7. Among our models, the ones that were trained on large quantities of pseudo-labeled ST data (TowerFull, SpireNoBlocks, and SpireFull) achieve far higher scores on direct ST than the one that did not (SpireNoPseudo). This indicates that merely performing CPT with ASR and MT data is not enough to achieve generalization to the task of direct ST, even if the model excels at both ASR and MT. Indeed, we also attempted direct ST with SpireBase and it failed to produce output in the target language, even when given few-shot prompts.
We also observe that SpireNoBlocks performs nearly as well at direct ST as SpireFull, even though its MT performance is much poorer (see Table 5 and 6), showing that, surprisingly, competence at MT is not very helpful for direct ST. SpireFull achieves the best self-cascaded performance by a significant margin for both datasets, outperforming both models that were trained with fewer speech samples (TowerFull and SpireNoPseudo), and the model that was not tuned on MT (SpireNoBlocks). This suggests that, unlike direct ST, both ASR and MT competencies are necessary for a strong self-cascade performance. Although SpireFull does not reach the direct ST performance of SeamlessM4T, it manages to achieve competitive performance despite using far less ST data Barrault et al. (2023). Our self-cascading experiments additionally demonstrate that SpireFull maintains greater robustness to its own outputs than SeamlessM4T.
@@ 5 Conclusion
In this work we presented Spire, a simple and effective recipe for adapting a text-based, translation-specialist LLM to the speech modality while preserving the original performance on text-based tasks. We investigated the impact of speech integration on two stages of LLM adaptation, CPT and IT, finding that both contribute to the final model’s performance on speech tasks. Our results demonstrate that we are able to successfully integrate a new modality without compromising the original model’s capabilities. Spire achieves competitive performance on ASR, while its MT abilities remain on par with the original Tower model. Finally, for the ST task, we find that the leveraging ASR and MT data does not directly transfer to ST performance. Nonetheless, the model achieves promising performance with both direct and self-cascaded ST.
As future work, we intend to extend this recipe to multilingual settings by replacing our English HuBERT speech component by the multilingual mHuBERT-147 Boito et al. (2024). We also plan to leverage the flexibility of DSU modeling to investigate the integration of speech generation tasks. To benefit the research community, we only use publicly available and licensed data to train our models, making our results reproducible.
@@
The downstream tasks we evaluate on are restricted to MT and ASR/ST, which provide an idea of the model performance but do not give us the full picture. We plan to address this by utilizing the LM-harness evaluation Gao et al. (2024) to evaluate on a suite of text-based benchmarks such as MMLU (Multitask Language Understanding) Hendrycks et al. (2021b, a), Arc (Commonsense Reasoning) Clark et al. (2018), Belebele (Reading Comprehension) Bandarkar et al. (2024), and HellaSwag (Sentence Completion) Zellers et al. (2019). Lastly, our model handles speech and text on the input side but is currently limited to generating only text.

### v3

@@ 4 Experiments
We evaluate our models across three tasks: ASR, MT, and ST. First, we present our results for ASR (§4.1), confirming the new capabilities Spire has in the speech domain. We then present MT results (§4.2), demonstrating that the speech performance does not come at the expense of the original model’s MT performance. Finally, we present results for ST (§4.3) to investigate model performance on a task that requires both ASR and MT capabilities.
Evaluation Setup
Across models and tasks, we perform inference with greedy decoding with a maximum of 256 generated tokens. For the Tower and Spire models, we decode with vllm Kwon et al. (2023). However, since vllm does not support all of our baselines, we use alternative libraries (transformers; Wolf et al., 2019) where necessary. Unless specified otherwise, we use zero-shot prompts for all models and tasks.
4.1 ASR
Datasets and Metrics
We evaluate ASR performance across multiple test sets, in order to cover a variety of recording styles: LibriSpeech (LS) test-clean and test-other Panayotov et al. (2015), FLEURS Conneau et al. (2023), and VoxPopuli.111111For CPT models, LS is an in-domain evaluation because its training set is part of MLS. We report the Word Error Rate (WER) between the hypotheses and gold transcripts, after Whisper normalization (Radford et al., 2023).
Baselines
We include the following models:
Whisper (Radford et al., 2023) is an encoder-decoder transformer trained on over 5 million hours of labeled data that performs multilingual ASR and to-English ST. We report results for Whisper-base (74M parameters) and Whisper-large-v3 (1.5B parameters).
SeamlessM4T (Barrault et al., 2023) is an encoder-decoder transformer trained on 406K hours of speech that performs ASR, ST and MT across 100 languages. We report results for SeamlessM4T-large-v2 (2.3B parameters).
SALMONN Tang et al. (2024) integrates a pre-trained text LLM with separate speech and audio encoders into a single multimodal model.121212SALMONN uses 4400 hours of speech/audio data in the IT phase but does not specify the large amount of pre-training ASR and audio captioning data used. SALMONN uses a LoRA adapter Hu et al. (2022) to align the spaces.
SpiritLM Nguyen et al. (2025) is a decoder-only model, trained from Llama-2 on 307B tokens of text, 458K hours of unlabeled speech, and 111K hours of labeled speech. As in Spire, it uses HuBERT DSUs.
HuBERT-large+CTC is a CTC-based ASR model trained using the same speech representation model we use for DSU generation, and using the same ASR data from the IT stage (Section 3.3.1).131313The hyperparameters are described in Appendix B. Unlike Spire, this model has access to a very powerful speech representation backbone. However, it lacks strong language modeling capabilities.
| LibriSpeech | FLEURS | VoxPopuli | ||
| Clean | Other | |||
| Whisper-base | 5.0 | 11.9 | 12.1 | 9.8 |
| Whisper-large-v3 | 1.8 | 3.7 | 5.8 | 9.2 |
| SeamlessM4T | 2.6 | 4.9 | 8.1 | 7.5 |
| SALMONN | 2.4 | 5.3 | 9.3 | 8.9 |
| Qwen2-Audio | 1.6 | 3.9 | 6.6 | 6.5 |
| SpiritLM | 6.0* | 11.0* | - | - |
| HuBERT-large+CTC | 4.3 | 7.6 | 11.4 | 14.7 |
| Our models | ||||
@@ 4.1 ASR
*We were unable to reproduce SpiritLM’s ASR performance; therefore, we report their self-reported LS results using ten-shot prompts.
@@ Results
Our results are presented in Table 2. SpireFull’s performance demonstrates that performing both the CPT and IT stages is an effective strategy to give speech capabilities to a text LLM. On the other hand, SpireBase does not consistently show reasonable speech performance; however, on FLEURS and VoxPopuli we obtain somewhat strong results in the zero-shot settings, which is surprising given that non-instruction-tuned models often struggle to work out-of-domain without in-context learning examples.141414We also tried prompting SpireBase with few-shot examples, but the results were much worse, possibly because the length of the DSU sequences led to in-context examples that were too long for the model to handle effectively.
Although SpireFull does not match the performance of SeamlessM4T, Whisper-large-v3, SALMONN, or Qwen2-Audio, these were trained on far more speech data than our models (around 10x for Qwen2-Audio and SeamlessM4T). Given this training data gap, it is notable that SpireFull does outperform Whisper-base on LS and FLEURS, and SpiritLM on all benchmarks SpiritLM reports, at a fraction of the speech data.
SpireFull also outperforms the HuBERT-large+CTC baseline on three out of four datasets. This is an impressive result given that the CTC model has access to continuous features, which SpireFull lacks. We believe this demonstrates that our compressed discrete representations capture the speech signal well enough to support speech-to-text tasks.
4.2 MT
@@
[…]
Having demonstrated that our training approach works well to initially equip Tower with speech processing capabilities, we now turn to MT to investigate whether Spire can maintain Tower’s strong performance on MT despite its speech-centric CPT.
Datasets and Metrics
We evaluate on two datasets for MT: FLORES-200 Team et al. (2024), which covers Spire’s languages, and the WMT23 test set (Kocmi et al., 2023), which covers en{de, ru, zh}.
We report COMET-22 (COMET; Rei et al., 2022a) and spBLEU151515nrefs:1|case:mixed|eff:no|tok:flores200|
smooth:exp|version:2.5.1 Papineni et al. (2002) scores via the SacreBLEU toolkit Post (2018).
Baselines
We compare the Spire models to the text-to-text translation performance of SeamlessM4T. Additionally, we report the performance of Towerbase-7b and Towerinstruct-7b.
Results
Our results show that even after the speech-centric CPT and mixed speech and text IT stage, the Spire models retain the original text-only performance of Tower on both FLORES (Table 3) and WMT23 (Table 5). This indicates that neither CPT nor IT on speech data negatively impacts the model’s ability to perform MT. This is true for both SpireBase, which achieves performance comparable to Towerbase; and for IT models, where SpireFull slightly surpasses the performance of Towerinstruct on enxx. Table 6 shows that between Towerinstruct and SpireFull, neither model consistently shows a significant improvement over the other. SpireFull also outperforms SeamlessM4T by both metrics on all WMT23 language pairs, and for both enxx and xxen on FLORES.
Translation-related Tasks
We follow the evaluation set-up from Tower (Alves et al., 2024) to additionally evaluate Spire on translation-related tasks. In Table 4 we report our results on APE for en{de, ru, zh} and NER for {de, en, es, fr, it, pt, zh}. Spire performs similarly to Towerinstruct across both tasks and all language directions, maintaining the original text-only capabilities even after training on speech data.
| APE | NER | ||
| enxx | xxen | Multilingual | |
| TowerInstruct-7b | 83.08 | 80.29 | 71.56 |
[…]
@@
| Corpus | Metric | TI Better | SF Better | NS |
| FLORES | spBLEU | 3 | 4 | 11 |
| COMET | 1 | 1 | 16 | |
| WMT23 | spBLEU | 0 | 2 | 4 |
| COMET | 2 | 2 | 2 |
4.3 ST
@@
As well as direct ST, we report self-cascades, in which each model transcribes the audio before translating its own output to the target language (i.e., ASR followed by MT).
Baselines
We compare Spire to SeamlessM4T in both direct and cascaded settings. We also report the results of SALMONN and Qwen2-Audio, which are both 7B parameter models, like Spire. However, SALMONN and Qwen2-Audio do not support text-to-text translation, so we use them only for direct ST.161616Although Whisper is frequently used for ST, we exclude it because it only supports to-English translation, whereas Spire is a from-English ST model. Therefore ST comparisons between the two models are impossible. There are also coverage differences between the models: while SeamlessM4T can handle all of Spire’s language pairs, neither SALMONN nor Qwen2-Audio supports enko; SALMONN also does not support enru.
@@
| ende | enzh | |||
| C22 | spB | C22 | spB | |
| Self-cascade | ||||
| SeamlessM4T | 72.40 | 21.7 | 72.32 | 17.0 |
[…]
| Direct | ||||
| SALMONN | 74.98 | 22.7 | 80.92 | 27.8 |
| Qwen2-Audio | 82.29 | 34.5 | 85.27 | 38.7 |
| SeamlessM4T | 85.95 | 42.3 | 83.62 | 31.3 |
[…]
Our FLEURS spBLEU ST results are reported in Table 8. For brevity, COMET-22 scores are reported in Appendix C. SeamlessM4T performs best at direct ST for all language pairs except enzh. Among the 7B parameter models, SpireFull is the best direct model on average, notably beating SALMONN on all language pairs except enzh. It also outperforms Qwen2-Audio on 6 out of 8 language pairs that Qwen2-Audio supports, and ties or beats it for all except enzh and ende.
Performance on CoVoST-2 (Table 7) tells a different story. Although SpireFull maintains its advantage over SeamlessM4T in self-cascaded translation, it attains the worst performance on enzh, while performing similarly to SALMONN for ende. This shows that the direct ST performance of SpireFull is dataset-dependent, which could be a consequence of its relatively small training data.
SpireFull achieves the best self-cascaded performance by a significant margin for both datasets, outperforming SeamlessM4T by a large margin in this setting. This demonstrates that SpireFull maintains greater robustness to its own outputs than SeamlessM4T, supporting the insight that LLM-based translation models can be very robust to perturbations (Peters and Martins, 2025).
| de | es | fr | it | ko | nl | pt | ru | zh | |||
| Self-Cascade | |||||||||||
| SeamlessM4T | 24.2 | 21.5 | 37.7 | 18.9 | 12.5 | 16.9 | 28.2 | 27.1 | 14.6 | 23.1 | 22.4 |
| SpireFull | 38.1 | 29.4 | 45.3 | 31.2 | 23.1 | 31.2 | 42.9 | 33.5 | 29.0 | 35.3 | 33.7 |
| Direct | |||||||||||
| SeamlessM4T | 39.2 | 28.0 | 48.1 | 30.6 | 21.5 | 30.8 | 47.5 | 34.3 | 23.2 | 35.3 | 33.7 |
| SALMONN | 25.5 | 20.8 | 34.3 | 16.7 | 0.1 | 20.5 | 32.6 | 3.1 | 21.9 | 24.6 | 19.5 |
| Qwen2-Audio | 31.8 | 23.5 | 31.3 | 23.5 | 5.4 | 22.3 | 36.1 | 23.7 | 24.7 | 27.6 | 24.7 |
[…]
5 Analysis
The key innovation of our approach is the application of the CPT followed by IT paradigm to discretized speech allowing us to build upon existing text-only capabilities of our base model. Here, we analyze how the composition of these two training phases contributes overall to model performance across all tasks previously evaluated. To that end, we consider several variants of SpireBase and SpireFull which are described in Table 9 and whose results are reported in Table 10.
i) No CPT was performed and IT was performed with the entire IT data mix (TowerFull);
ii) CPT was performed and no data from TowerBlocks was seen during IT (SpireNoBlocks), and
iii) CPT was performed and pseudo-labeled ST data and FLEURS were omitted from the IT data mix (SpireNoPseudo).
We report additional datasets in Appendix D.
| Model | Base Model | CPT | IT | |||
| Speech | Text | Speech | Pseudo | Text | ||
| TowerFull | TowerBase-7B | ✗ | ✗ | ✓ | ✓ | ✓ |
[…]
| Spire Variants | ||||||
| SpireNoBlocks | SpireBase | ✓ | ✓ | ✓ | ✓ | ✗ |
| SpireNoPseudo | SpireBase | ✓ | ✓ | ✓ | ✗ | ✓ |
Effectiveness of CPT and IT
Our previous results demonstrated that using both CPT and IT was the most effective strategy. The performance gap between SpireFull and the TowerFull on ASR (5.3 points in LS test-clean) further shows that IT alone is also not as effective. However, for ST we observe that performing only IT leads to a strong model that is capable of performing speech translation. This contrasts from SpireBase, for which we also attempted direct ST but the model failed to produce output in the target language, even when given few-shot prompts. Despite the impressive results from TowerFull, we still observe the best performance by SpireFull showing that while the effect of CPT is not as drastic as in the case of ASR, we still observe gains with a speech-centric CPT phase.
Modality Interplay
Our results show that text and speech modalities are orthogonal to each other. Specifically, the performances of TowerFull and SpireFull show that speech-centric CPT does not degrade the text performance of the base model. However, MT quality suffers when TowerBlocks is removed from the IT data, as is shown by SpireNoBlocks’s much weaker performance than SpireFull. Simultaneously, SpireFull performs on par with SpireNoBlocks on both ASR and ST, indicating that adding text instructions also does not degrade performance on speech tasks. It is worth highlighting that a model strong at both MT and ASR (SpireNoPseudo) does not lead to a strong ST model, showing surprisingly that competence at MT is not very helpful for direct ST.
| ASR | MT | ST | |||||
| enxx | xxen | enxx | |||||
| WER | C22 | spB | C22 | spB | C22 | spB | |
[…]
| TowerFull | 9.5 | 88.57 | 39.4 | 88.17 | 41.7 | 79.10 | 26.1 |
| SpireNoBlocks | 4.1 | 82.98 | 34.2 | 85.93 | 36.1 | 81.11 | 27.1 |
| SpireNoPseudo | 3.9 | 88.40 | 38.9 | 88.22 | 42.0 | 62.80 | 27.1 |
@@ 6 Conclusion
In this work we presented Spire, a simple and effective recipe for adapting a text-based, translation-specialist LLM to the speech modality while preserving the original performance on text-based tasks. We investigated the impact of speech integration on two stages of LLM adaptation, CPT and IT, finding that both contribute to the final model’s performance on speech tasks. Our results demonstrate that we are able to successfully integrate a new modality without compromising the original model’s capabilities. Spire achieves competitive performance on ASR, while its MT abilities remain on par with the original Tower model. Finally, for the ST task, we find that the leveraging ASR and MT data does not directly transfer to ST performance. Nonetheless, the model achieves promising performance with both direct and self-cascaded ST. To benefit the community, we only used publicly available and licensed data to train our models, making our results reproducible. As future work, we intend to extend this recipe to multilingual settings by replacing our English HuBERT speech component by the multilingual mHuBERT-147 Boito et al. (2024).
@@
The downstream tasks we evaluate on are restricted to MT and ASR/ST, which provides an idea of the model performance but do not give us the full picture. We plan to address this by utilizing the LM-harness evaluation Gao et al. (2024) to evaluate on a suite of text-based benchmarks such as MMLU (multitask language understanding) Hendrycks et al. (2021a, b), Arc (commonsense reasoning) Clark et al. (2018), Belebele (reading comprehension) Bandarkar et al. (2024), and HellaSwag (sentence completion) Zellers et al. (2019). Lastly, our model handles speech and text on the input side but is currently limited to generating only text.